# Throwing Vines at the Wall: Structure Learning via Random Search

### v1

@@ Abstract
Vine copulas offer flexible multivariate dependence modeling and have become widely used in machine learning, yet structure learning remains a key challenge. Early heuristics like the greedy algorithm of Dissmann are still considered the gold standard, but often suboptimal. We propose random search algorithms that improve structure selection and a statistical framework based on model confidence sets, which provides theoretical guarantees on selection probabilities and a powerful foundation for ensembling. Empirical results on several real-world data sets show that our methods consistently outperform state-of-the-art approaches.
@@ 1 INTRODUCTION
Copulas provide a flexible framework for modeling complex multivariate distributions by disentangling marginal behavior from the dependence structure (see, e.g., Nelsen,, 2007; Joe,, 2014). Among them, the class of vine copulas (Bedford and Cooke,, 2001; Aas et al.,, 2009) has become particularly popular in machine learning because it balances flexibility with tractability and scales better than generic high-dimensional copulas. Recent applications span a wide range of areas, including domain adaptation (Lopez-Paz et al.,, 2013), variational inference (Tran et al.,, 2015), causal inference (Lopez-Paz,, 2016), clustering (Tekumalla et al.,, 2017), Bayesian optimization (Park et al.,, 2024), conformal prediction (Park et al.,, 2025), and generative modeling (Kulkarni et al.,, 2018; Konstantelos et al.,, 2018; Sun et al.,, 2019; Tagasovska, Ackerer, and Vatter,, 2019).
A vine copula consists of two components: a vine structure, an undirected graph composed of nested trees, and bivariate (conditional) copulas, called pair-copulas, associated with its edges. The number of possible vine structures on $d$ variables equals $2^{(d-3)(d-2)/2-1}d!$ (Morales-Nápoles,, 2011; Joe et al.,, 2011). This super-exponential growth is of the same order as that of directed acyclic graphs with the dimension, making exhaustive search infeasible beyond a few variables. As a result, nearly all applications rely on heuristics such as the greedy spanning tree algorithm of Dissmann et al., (2013) or the forward selection procedure of Kraus and Czado, 2017a .
Despite their simplicity and lack of theoretical foundation, these heuristics remain the default: numerous attempts at improving on them have met with limited success (see Section 2.4). The recent review of Czado and Nagler, (2022) concludes that such heuristics still represent the state of the art, and that effective structure selection remains an important open challenge in vine methodology.
Figure 1: Average relative improvement in out-of-sample log-likelihood over Dissmann et al., (2013).
In Figure 1, we illustrate the potential for improvement by comparing the out-of-sample log-likelihood of Dissmann et al., (2013) to that of vines sampled uniformly at random, selecting the best candidate using an independent validation set (see Section 4 for details). From the picture, it is clear that the greedy heuristic is suboptimal, and that random search performance improves with the number of candidates.
@@
  • We integrate random search with a state-of-the-art algorithm for constructing model confidence sets (Kim and Ramdas,, 2025) tailored to vine structures, prove its validity in our context, and provide an efficient implementation. This enables in-sample assessment of whether alternatives outperform a baseline heuristic and yields a principled set of competitive models.
  • We demonstrate that vine regression methods based on averaging over the model confidence set consistently outperform single-model approaches, including the current state of the art.
The new algorithms are conceptually simple and straightforward to implement. Yet, they represent arguably the first substantive improvement in vine structure selection since the introduction of the original heuristics. They provide immediate benefits for applications of vine copulas in machine learning and related fields, unlocking additional performance gains.
@@
The remainder of this paper is organized as follows. In Section 2, we review the necessary background on vine copulas and structure learning. In Section 3, we introduce our new structure learning algorithms. In Section 4, we present an extensive empirical evaluation of our methods and compare them to state-of-the-art approaches. Finally, we conclude in Section 6.
@@ 2.1 Copulas
Mathematically, a copula is a multivariate distribution with standard uniform margins. From Sklar, (1959), a random vector $\bm{X}\in\mathds{R}^{d}$ has joint distribution $F$ with univariate marginal distributions $F_{1},\dots,F_{d}$, if and only if there exists a copula $C$ such that
$F\left(x_{1},\,\ldots,\,x_{d}\right)=C\left\{F_{1}(x_{1}),\,\ldots,\,F_{d}(x_{d})\right\}\quad\forall\,x\in\mathds{R}^{d},$
@@
Taking logarithm on both sides further implies that the joint log-likelihood is the sum of the marginal and copula log-likelihoods, which can be exploited in a two-step estimation procedure (Genest et al.,, 1995; Joe and Xu,, 1996): first estimate each of the marginal distributions, and then the copula based on the pseudo-observations $\widehat{U}_{i}=\widehat{F}_{i}(X_{i})$, $i=1,\dots,d$.
@@ 2.2 Vines
Following the seminal work of Joe, (1996) and Bedford and Cooke, (2001, 2002), any copula density $c$ can be decomposed into a product of d(d1)/2d(d-1)/2 bivariate (conditional) copula densities. The order of conditioning in this decomposition can be organized using the following graphical structure.
@@
An example of a vine tree sequence is shown in Figure 2.
12341,22,33,41,3||22,4||31,22,33,41,3||22,4||31,4||2,3
Figure 2: A vine structure on four variables.
A vine copula identifies each edge $e$ of the vine structure, with a label $\{j_{e},k_{e}|D_{e}\}$ and a bivariate copula $C_{j_{e},k_{e}|D_{e}}$, called a pair-copula. The sets $\{j_{e},k_{e}\}\subset\{1,\ldots,d\}$ and $D_{e}\subset\{1,\ldots,d\}\setminus\{j_{e},k_{e}\}$ are called the conditioned set and conditioning set, respectively, and we refer to Czado and Nagler, (2022) for a precise definition. As for $C_{j_{e},k_{e}|D_{e}}$, it describes the dependence between $U_{j_{e}}$ and $U_{k_{e}}$ conditionally on the subvector $\bm{U}_{D_{e}}$ corresponding to the conditioning set $D_{e}$. The corresponding copula density for $\bm{u}\in[0,1]^{d}$ then factorizes as
$c(\bm{u})$ $=\prod_{t=1}^{d-1}\prod_{e\in E_{t}}c_{j_{e},k_{e}|D_{e}}\bigl(u_{j_{e}|D_{e}},u_{k_{e}|D_{e}}|\bm{u}_{D_{e}}\bigr),$ (1)
@@
Thanks to the proximity condition, $u_{j_{e}|D_{e}}$ for any edge $e\in E_{t+1}$ can be computed recursively from the pair-copulas in trees $1,\dots,t$. This way, the parameters of vine copulas can be estimated by iterating between parameter estimation and conditional distribution evaluations, going from the first tree to the last (Aas et al.,, 2009). A similar decomposition holds when some variables are discrete, effectively by replacing derivatives of the distribution function by finite differences (Panagiotelis et al.,, 2012; Funk et al.,, 2025).
@@ 2.3 Why the Structure Matters
Equation (1) provides a general decomposition of any copula density, regardless of the vine structure. In practice, however, the conditional copula densities involved are high-dimensional and intractable. The key to the practical success of vine copulas is the simplifying assumption that these conditional copulas do not depend on the conditioning value, i.e.,
$c(\bm{u})$ $=\prod_{t=1}^{d-1}\prod_{e\in E_{t}}c_{j_{e},k_{e}|D_{e}}\bigl(u_{j_{e}|D_{e}},u_{k_{e}|D_{e}}\bigr).$ (2)
This reduces the model to a collection of bivariate copulas, which are straightforward to estimate, yet still highly flexible. Each pair-copula can be chosen independently, allowing complex dependence patterns with varying strength, asymmetry, and non-linear tail behavior. More in-depth discussions of the assumption can be found in Stoeber et al., (2013); Spanhel and Kurz, (2019); Nagler, (2025).
@@ 2.4 Related Work
The standard approach was introduced by Dissmann et al., (2013), whose greedy algorithm constructs a maximum spanning tree based on the absolute value of Kendall’s $\tau$. Variants using information criteria or goodness-of-fit $p$-values (Czado et al.,, 2013) offer no clear benefits but incur higher computational cost. Attempts to incorporate tests of the simplifying assumption (Kraus and Czado, 2017b, ) also failed to consistently improve performance. More elaborate search strategies based on MCMC (Gruber and Czado,, 2015, 2018) or neural networks (Sun et al.,, 2019) can yield gains but are prohibitively expensive, limiting their use to low dimensions. Randomized search variants have also been explored, such as sampling from common variable orders (Zhu et al.,, 2020) or Monte Carlo tree search with lookahead (Chang et al.,, 2019), but both approaches again scale poorly with dimension. Overall, the greedy heuristic of Dissmann et al., (2013) remains the default in practice (Czado and Nagler,, 2022).
@@
Specialized algorithms have been developed for regression tasks. These methods restrict attention to structures that admit closed-form conditional distributions, often star- or path-shaped trees (D- and C-vines). Some extend Dissmann’s approach (Chang and Joe,, 2019; Chang et al.,, 2019; Cooke et al.,, 2019), while others use forward variable selection with early stopping (Kraus and Czado, 2017a, ; Schallhorn et al.,, 2017; Zhu et al.,, 2021; Tepegjozova et al.,, 2022). Their applicability is thus narrower than general-purpose methods.
@@
Another line of work focuses on sparsity, i.e., identifying independence copulas among pairs. Approaches include selecting truncation or thresholding levels (Kurowicka,, 2011; Brechmann et al.,, 2012; Brechmann and Joe,, 2015; Joe,, 2018; Nagler et al.,, 2019), typically combined with spanning-tree heuristics. A different perspective exploits connections between vine copulas and Gaussian DAGs (Müller and Czado,, 2018; Müller and Czado, 2019a, ; Müller and Czado, 2019b, ), aiming primarily to detect conditional independencies rather than fully optimizing the vine structure.
@@ 3.1 Setup
Given a vine structure $\mathcal{V}$ and a dataset $\mathcal{D}$, we assume access to an algorithm that fits an estimated joint density $\widehat{f}_{\mathcal{V},\mathcal{D}}\in\mathcal{F}$, with $\mathcal{F}$ the space of probability densities in $\mathds{R}^{d}$. This density is composed of estimated marginal distributions and a vine copula density of the form (2) with structure $\mathcal{V}$.
To evaluate different models, let $L\colon\mathcal{F}\times\mathds{R}^{d}\to\mathds{R}$ denote a loss functional, where $L(f,\bm{Z})$ quantifies the quality of $f\in\mathcal{F}$ based on an observation $\bm{Z}$. Common choices include the negative log-likelihood $L(f,\bm{Z})=-\log f(\bm{Z})$ for generative modeling, or similarly $L(f,(\bm{X},Y))=-\log f(Y|\bm{X})$ for regression-type problems, and we refer to Gneiting and Katzfuss, (2014) for a comprehensive review in the context of probabilistic forecasting.
The expected loss is $L(f)=\mathds{E}_{\bm{Z}}[L(f,\bm{Z})]$, and the goal of model selection is to find, among a set of candidates, the one with minimal expected loss. In practice however, $L(f)$ is unknown and must be estimated from data, which can be done, e.g., via hold-out validation or cross-validation (Stone,, 1974).
@@ 3.2 Random Search Algorithm
Algorithm 1 Hold-out Random Search for Vine Structure Learning.
@@
2:Optimal vine structure $\widehat{\mathcal{V}}_{RS}$.
@@
4:Generate candidate vine structures $\mathcal{V}_{1},\dots,\mathcal{V}_{M}$ uniformly at random from the set of all structures using the algorithm of Joe et al., (2011).
5:Fit the candidates on the training part of the data, yielding estimated densities $\widehat{f}_{\mathcal{V}_{1},\mathcal{D}_{\text{train}}},\dots,\widehat{f}_{\mathcal{V}_{M},\mathcal{D}_{\text{train}}}$.
6:For each model, compute the average validation loss
$L_{\text{val}}(\mathcal{V}_{k})=\frac{1}{|\mathcal{D}_{\text{val}}|}\sum_{\bm{x}\in\mathcal{D}_{\text{val}}}L(\widehat{f}_{\mathcal{V}_{k},\mathcal{D}_{\text{train}}},\bm{x}).$
7:Select the model with minimal validation loss:
$\widehat{\mathcal{V}}_{RS}=\operatorname*{arg\,min}L_{\text{val}}(\mathcal{V}_{k}).$
@@
To identify a suitable vine structure, we propose a simple and easy-to-implement algorithm based on hold-out random search, summarized in Algorithm 1. The procedure consists of splitting the data into training and validation parts, generating candidate structures at random, fitting them on the training data, and selecting the one with the smallest validation loss.
@@
For a typical vine model $\widehat{f}_{\mathcal{V},\mathcal{D}}$, both training and inference incurs an $O(d^{2})$ cost per observation, where the factor $d^{2}$ comes from the quadratic number of pair-copulas in the model. This leads to an overall runtime complexity of $O(Mnd^{2})$ for Algorithm 1. Note that it is embarrassingly parallel over the $M$ candidates, as well as, for a single candidate, tree-wise over pair-copulas. Furthermore, one can reduce the dependence on $d$ from quadratic to linear by truncating the vine at a fixed level (Kurowicka,, 2011; Brechmann et al.,, 2012; Brechmann and Joe,, 2015; Joe,, 2018; Nagler et al.,, 2019).
@@
While the approach naturally extends to cross-validation, we argue that the additional computational budget is better invested in evaluating more random candidates; see Figure 1 and Section 4. In addition, cross-validation does not solve the problem that the selected model may be statistically indistinguishable from other candidates, which we address in the next subsection.
3.3 Vine Structure Confidence Sets

### v2

@@ Abstract
Vine copulas offer flexible multivariate dependence modeling and have become widely used in machine learning. Yet, structure learning remains a key challenge. Early heuristics, such as Dissmann’s greedy algorithm, are still considered the gold standard but are often suboptimal. We propose random search algorithms and a statistical framework based on model confidence sets, to improve structure selection, provide theoretical guarantees on selection probabilities, and serve as a foundation for ensembling. Empirical results on real-world data sets show that our methods consistently outperform state-of-the-art approaches.
@@ 1 Introduction
Copulas provide a flexible framework for modeling complex multivariate distributions by disentangling marginal behavior from the dependence structure (see, e.g., nelsen2007introduction; joe2014). Among them, the class of vine copulas (bedfordcooke2001; aasczadofrigessibakken2009) has become particularly popular in machine learning because it balances flexibility with tractability better than more classical copula families. Recent applications span a wide range of areas, including domain adaptation (Lopez-Paz2013), variational inference (Tran2015), causal inference (Lopez-Paz2016), clustering (tekumalla2017vine), Bayesian optimization (botied), conformal prediction (park2025semiparametric), and generative modeling (kulkarni2018nonparametric; Konstantelos2018; sun2019; Tagasovska2018b).
A vine copula on $d$ variables consists of two components: its vine structure, a nested sequence of undirected trees, and d(d1)/2d(d-1)/2 bivariate (conditional) copulas, called pair-copulas, indexed by its edges. Vine copulas are most popular on problems from a handful to a few dozen variables. In this regime, the number of pair-copulas remains tractable, while the flexibility of the model can capture complex dependence patterns. However, there exists $2^{(d-3)(d-2)/2-1}d!$ possible vines on $d$ variables (morales2011count; joe2011regular), which is of the same order as the number of directed acyclic graphs on $d$ nodes. This super-exponential growth makes exhaustive search infeasible beyond a few variables. As a result, nearly all applications rely on greedy heuristics such as the spanning tree algorithm of Dissmann2013 or the forward selection procedure of kraus2017d.
Despite their simplicity and lack of theoretical foundation, numerous attempts at improving on them had very limited success (see Section˜2.4). The recent review of Czado2022 concludes that such heuristics still represent the state of the art, and that effective structure selection remains an important open challenge in vine methodology.
Figure 1: Average improvement in out-of-sample negative log-likelihood over Dissmann2013.
@@
  • We integrate random search with a state-of-the-art algorithm for constructing model confidence sets (kim2025locally) tailored to vine structures, prove its validity in our context, and provide an efficient implementation. This enables in-sample assessment of whether alternatives outperform a baseline heuristic and yields a principled set of competitive models.
  • We demonstrate that vine regression methods based on averaging over the model confidence set consistently outperform single-vine approaches, including the current state of the art.
In Figure˜1, we illustrate the potential for improvement by comparing the out-of-sample negative log-likelihood of Dissmann2013 to that of vines sampled uniformly at random, selecting the best candidates using an independent validation set (see Section˜4 for details). From the picture, it is clear that the greedy heuristic is suboptimal, and that random search performance improves with the number of candidates.
The new algorithms are conceptually simple, straightforward to implement, and provide immediate benefits for applications of vine copulas in machine learning and related fields.
@@
In Section˜2, we review the necessary background on vine copulas and structure learning. In Section˜3, we introduce our new structure learning algorithms. In Section˜4, we present an extensive empirical evaluation of our methods and compare them to state-of-the-art approaches. Finally, we conclude in Section˜5.
@@ 2.1 Copulas
Mathematically, a copula is a multivariate distribution with standard uniform margins. From sklar1959, a random vector $\bm{X}\in\mathds{R}^{d}$ has joint distribution $F$ with univariate marginal distributions $F_{1},\dots,F_{d}$, if and only if there exists a copula $C$ such that, for each $x\in\mathds{R}^{d}$,
$F\left(x_{1},\,\ldots,\,x_{d}\right)=C\left\{F_{1}(x_{1}),\,\ldots,\,F_{d}(x_{d})\right\},$
@@
Taking logarithm on both sides further implies that the joint log-likelihood is the sum of the marginal and copula log-likelihoods, which can be exploited in a two-step estimation procedure (genest1995semiparametric; joe1996estimation): first estimate each of the marginal distributions, and then the copula based on the pseudo-observations $\widehat{U}_{i}=\widehat{F}_{i}(X_{i})$, $i=1,\dots,d$.
@@ 2.2 Vines
Following the seminal work of joe1996 and bedfordcooke2001; Bedford2002, any copula density $c$ can be decomposed into a product of d(d1)/2d(d-1)/2 bivariate (conditional) copula densities. The order of conditioning in this decomposition can be organized using the following graphical structure.
12341,22,33,41,3||22,4||31,22,33,41,3||22,4||31,4||2,3
Figure 2: A vine structure on four variables.
@@
A simple example of a vine tree sequence is shown in Figure˜2. A vine copula identifies each edge $e$ of the vine structure, with a label $\{j_{e},k_{e}|D_{e}\}$ and a bivariate copula $C_{j_{e},k_{e}|D_{e}}$, called a pair-copula. The sets $\{j_{e},k_{e}\}\subset\{1,\ldots,d\}$ and $D_{e}\subset\{1,\ldots,d\}\setminus\{j_{e},k_{e}\}$ are called the conditioned set and conditioning set, respectively, and we refer to Czado2022 for a precise definition. As for $C_{j_{e},k_{e}|D_{e}}$, it describes the dependence between $U_{j_{e}}$ and $U_{k_{e}}$ conditionally on the subvector $\bm{U}_{D_{e}}$ corresponding to the conditioning set $D_{e}$. The corresponding copula density for $\bm{u}\in[0,1]^{d}$ then factorizes as
$c(\bm{u})$ $=\prod_{t=1}^{d-1}\prod_{e\in E_{t}}c_{j_{e},k_{e}|D_{e}}\bigl(u_{j_{e}|D_{e}},u_{k_{e}|D_{e}}\mid\bm{u}_{D_{e}}\bigr),$ (1)
@@
Thanks to the proximity condition, $u_{j_{e}|D_{e}}$ for any edge $e\in E_{t+1}$ can be computed recursively from the pair-copulas in trees $1,\dots,t$. This way, the parameters of vine copulas can be estimated by iterating between parameter estimation and conditional distribution evaluations, going from the first tree to the last (aasczadofrigessibakken2009). A similar decomposition holds when some variables are discrete, effectively by replacing derivatives of the distribution function by finite differences (panagiotelis2012; funk2025towards).
@@ 2.3 Why the Structure Matters
Equation (1) provides a general decomposition of any copula density, regardless of the vine structure. The key to the practical success of vine copulas is a simplifying assumption that the conditional copulas in (1) do not depend on the conditioning value $\bm{u}_{D_{e}}$. This reduces the model to a collection of bivariate copulas, which are straightforward to estimate, yet still highly flexible. Each pair-copula can be chosen independently, allowing complex dependence patterns with varying strength, asymmetry, and non-linear tail behavior. More in-depth discussions of the assumption can be found in stoeber2013simplified; spanhel2019simplified; nagler2025simplified.
@@ 2.4 Related Work
The standard approach was introduced by Dissmann2013, whose greedy algorithm constructs a maximum spanning tree based on the absolute value of Kendall’s $\tau$. Variants using information criteria or goodness-of-fit $p$-values (czado2013selection) offer no clear benefits but incur higher computational cost. Attempts to incorporate tests of the simplifying assumption (Kraus2017) also failed to consistently improve performance. More elaborate search strategies based on MCMC (gruber2015sequential; gruber2018) or neural networks (sun2019) can yield gains but are prohibitively expensive: even in the low-dimensional examples provided in the papers, the algorithm run times are on the order, or larger than, those of exhaustive search. Randomized search variants have also been explored. Sampling from common variable orders (zhu2020common) does not improve out-of-sample performance on real data and the Monte Carlo tree search approach with lookahead of chang19a only applies to truncated Gaussian models. Overall, the greedy heuristic of Dissmann2013 remains the default in practice (Czado2022).
@@
Specialized algorithms have been developed for regression tasks. These methods restrict attention to structures that admit closed-form conditional distributions, often star- or path-shaped trees (D- and C-vines). Some extend Dissmann’s approach (chang2019prediction; chang19a; cooke2019vine), while others use forward variable selection with early stopping (kraus2017d; Schallhorn2017; zhu2021simplified; tepegjozova2022nonparametric). Their applicability is thus narrower than general-purpose methods.
@@
Another line of work focuses on sparsity, i.e., identifying independence copulas among pairs, which becomes necessary in higher dimensions to keep models tractable. Approaches include selecting truncation or thresholding levels (kurowicka2011optimal; BrechmannCzadoAas2012; brechmann2015truncation; joe2018parsimonious; nagler2019model), typically combined with spanning-tree heuristics. A different perspective exploits connections between vine copulas and Gaussian DAGs (mueller2018representing; muller2019dependence; muller2019selection), aiming primarily to detect conditional independencies rather than fully optimizing the vine structure. These questions are important but somewhat orthogonal to the focus of this paper, which is the structure itself, rather than the pair-copulas.
@@ 3.1 Setup
Given a vine structure $\mathcal{V}$ and a dataset $\mathcal{D}$, we assume access to an algorithm that fits an estimated joint density $\widehat{f}_{\mathcal{V},\mathcal{D}}\in\mathcal{F}$, with $\mathcal{F}$ the space of probability densities in $\mathds{R}^{d}$. This density is composed of estimated marginal distributions and a simplified vine copula with structure $\mathcal{V}$.
To evaluate different models, let $L\colon\mathcal{F}\times\mathds{R}^{d}\to\mathds{R}$ denote a loss functional, where $L(f,\bm{Z})$ quantifies the quality of $f\in\mathcal{F}$ based on an observation $\bm{Z}$. Common choices include the negative log-likelihood $L(f,\bm{Z})=-\log f(\bm{Z})$ for generative modeling, or similarly $L(f,(\bm{X},Y))=-\log f(Y|\bm{X})$ for regression-type problems, and we refer to gneiting2014probabilistic for a comprehensive review in the context of probabilistic forecasting.
The expected loss is $L(f)=\mathds{E}_{\bm{Z}}[L(f,\bm{Z})]$, and the goal of model selection is to find, among a set of candidates, the one with minimal expected loss. In practice however, $L(f)$ is unknown and must be estimated from data, which can be done, e.g., via hold-out validation or cross-validation (stone1974cross).
@@ 3.2 Random Search Algorithm
Algorithm 1 Hold-out Random Search for Vines.
@@
2:Optimal vine structure $\widehat{\mathcal{V}}$.
@@
4:Generate candidate structures $\Theta=\{\mathcal{V}_{1},\dots,\mathcal{V}_{M}\}$.
5:For each candidate $\mathcal{V}\in\Theta$:
6:   Use the training set to fit a density $\widehat{f}_{\mathcal{V},\mathcal{D}_{\text{train}}}$.
7:   Compute the validation loss
$L_{\text{val}}(\mathcal{V})=\frac{1}{|\mathcal{D}_{\text{val}}|}\sum_{\bm{z}\in\mathcal{D}_{\text{val}}}L(\widehat{f}_{\mathcal{V},\mathcal{D}_{\text{train}}},\bm{z}).$
8:Select the structure with minimal validation loss:
$\widehat{\mathcal{V}}=\operatorname*{arg\,min}_{\mathcal{V}\in\Theta}L_{\text{val}}(\mathcal{V}).$
@@
To identify a suitable vine structure, we propose a simple and easy-to-implement algorithm based on hold-out random search, summarized in Algorithm˜1. The procedure consists of splitting the data into training and validation sets, generating candidate structures at random, fitting them on the training data, and selecting the one with the smallest validation loss.
@@
For a typical vine model $\widehat{f}_{\mathcal{V},\mathcal{D}}$, both training and inference incurs an $O(d^{2})$ cost per observation, where the factor $d^{2}$ comes from the quadratic number of pair-copulas in the model. This leads to an overall runtime complexity of $O(Mnd^{2})$ for Algorithm˜1. Note that it is embarrassingly parallel over the $M$ candidates, as well as, for a single candidate, tree-wise over pair-copulas. Furthermore, one can reduce the dependence on $d$ from quadratic to linear by truncating the vine at a fixed level (e.g., BrechmannCzadoAas2012).
For generating vine structures, we propose to sample uniformly at random using the algorithm of joe2011regular; see also joe2014 and its implementation in the software libraries VineCopula and pyvinecopulib (vinecopula2025; pyvinecopulib). Other sampling schemes, potentially involving the training data, could be considered as well and also fit the theoretical framework and results ahead; see Section˜D.3 for a sampling scheme based on local perturbations of the Dissmann2013 structure.
@@
While the approach naturally extends to cross-validation, we argue that the additional computational budget is better invested in evaluating more random candidates; see Figure˜1 and Section˜4. In addition, cross-validation does not solve the problem that the selected model may be statistically indistinguishable from other candidates, which we address in the next subsection.
3.3 Vine Confidence Sets